OpenStudy (anonymous):

What is the length of one side of a regular octagon if its perimeter is 72cm no answers @awillie98,@iihoop3 and others - will medal a maybe fan

OpenStudy (anonymous):

Im not eve sure how to do this

OpenStudy (anonymous):

lol its ok girl! no worries @TheSmartOne

OpenStudy (mathstudent55):

How many sides does an octagon have?

OpenStudy (anonymous):

8 of course

OpenStudy (anonymous):

9

OpenStudy (anonymous):

is the answer

OpenStudy (mathstudent55):

Great. What does "regular" mean in "regular octagon"?

OpenStudy (anonymous):

9*8 sides of the octagon is 72

OpenStudy (anonymous):

a regular octagon is just the regular kind ..wait I thought a octagon had 5 sides? sorry im just so tired

OpenStudy (anonymous):

lmao no it has 8 sides

OpenStudy (anonymous):

What is the measure of BAC ? A.28 B.83 C.55 D.93 sorry I just tired ...I wa thinking C

OpenStudy (mathstudent55):

No. You were correct the first time. An octagon has 8 sides. A regular polygon means that all sides are congruent, and all angles are congruent. Since it has 8 sides of the same length, and the perimeter is 72 cm, then you can divide (72 cm)/8 = 9 cm, so you get that each side is 9 cm long. If you were not told it is a regular octagon, you wouldn't be able to answer the question because the sides would have different lengths.

OpenStudy (mathstudent55):

What do you know about the sum of the measures of the angles of a triangle?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

@mathstudent55 can you help me out also

OpenStudy (anonymous):

can u tll me the answer I have a time limit and I don't want to get in trouble

OpenStudy (mathstudent55):

The sum of the measures of the angles of a triangle is 180 degrees. All three angles must add to 180. You were given the measures of two angles. Add them up. Then subtract the sum from 180, and that is the measure of the third angle.

OpenStudy (mathstudent55):

We'll do it quickly. What is 62 + 35 = ?

OpenStudy (anonymous):

97

OpenStudy (mathstudent55):

Great. 62 + 35 = 97 Now subtract the sum 97, from 180. What is 180 - 97 = ?

OpenStudy (anonymous):

83

OpenStudy (mathstudent55):

Correct. To double check, just add: Is 62 + 35 + 83 = 180? Yes, it is, so 83 is correct.

OpenStudy (mathstudent55):

Ok, gtg. I'll be back in a few hours.

OpenStudy (anonymous):

ok thx so much!

OpenStudy (mathstudent55):

You're welcome. If you see me on OS again and have any questions on math, just tag me or message me,.

OpenStudy (mathstudent55):

Bye.

OpenStudy (anonymous):

ok bye